OpenStudy (anonymous):

optimization: The sum of two nonnegative numbers is 36. Find the numbers if a. the difference of their square roots is to be as large as possible. b. the sum of their square roots is to be as large as possible.

OpenStudy (anonymous):

first i know to find the relationsihps so x + y = 36 so x = 36-y

OpenStudy (anonymous):

part b seems like it's the easiest to do, so i started with that one. so to find the optimal i take the derivative of root of x + root of y and set it equal to 0

OpenStudy (kainui):

So far so good.

OpenStudy (anonymous):

Right, you have \(x\) and \(36-x\) as your numbers. You also know \(x\geq 0\).

OpenStudy (anonymous):

alright so after substiuting in 36 - y for x in the equation i got: \[\sqrt{36 - y} + \sqrt{y}\]

OpenStudy (anonymous):

For part a, we have: \[ f(x) = \sqrt{x}-\sqrt{36-x} \]And you are trying to find the max.

OpenStudy (anonymous):

oh okay, we'll start with part a.

OpenStudy (anonymous):

Can you find the critical numbers?

OpenStudy (anonymous):

let's see i take the derivative set it equal to 0 and solve for y right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so for the derivative i got \[\frac{ 1 }{ 2\sqrt{x} } + \frac{ 1 }{ 2\sqrt{36-x} } = 0\]

OpenStudy (anonymous):

You want to multiply both sides by the LCM, which is \(2\sqrt x \sqrt{36-x}\)

OpenStudy (anonymous):

that means that i get \[\sqrt{36-x} + \sqrt{x} = 0\]

OpenStudy (anonymous):

now multiply by the conjugate.

OpenStudy (anonymous):

Actually, it's a bit weird... both of them are technically positive so...

OpenStudy (anonymous):

I think you forgot a negative in your derivative.

OpenStudy (anonymous):

Nevermind, your derivative is correct.

OpenStudy (solomonzelman):

find the common denominator, and solve for x.

OpenStudy (anonymous):

solving for y with the conjugates got me x = 18 and y = 18 which is the answer for part b

OpenStudy (solomonzelman):

let me check...

OpenStudy (anonymous):

Actually, I don't think \(x=18\) is a critical number.

OpenStudy (solomonzelman):

\(\LARGE\color{black}{\frac{1}{2\sqrt{x}}+\frac{1}{2\sqrt{36-x}}=0 }\) \(\LARGE\color{black}{\sqrt{x}\sqrt{36-x}=0 }\) critical numbers 0 or 36.

OpenStudy (solomonzelman):

(I have a prayer right now, excuse me... will see you:))

OpenStudy (anonymous):

okay no problem.

OpenStudy (anonymous):

The only critical numbers are \(x=0\) and \(x=36\), since it is undefined at these points. Also, since they're the interval boundaries.

OpenStudy (anonymous):

so then i just plug them into the equation and see what it is right?

OpenStudy (anonymous):

I 've got to go, but i think i get this. Thanks